# Contrastive Alignment with Semantic Gap-Aware Corrections in Text-Video Retrieval

### v4

@@ 3.2 Gap-Aware Increment Modeling via Pair-Specific $\Delta_{ij}$
$\Delta_{ij}^{(t)}=\psi\left(\eta(v_{j}-t_{i}),c;\Theta^{(t)}\right),$ (9)
where $\eta\in\{-1,+1\}$ determines the direction of the semantic gap depending on which modality receives the increment. The function $\psi$ is implemented as a single-layer cross-attention module, where the semantic difference $\eta(v_{j}-t_{i})$ serves as the query, and the context vector $c$—defined based on dataset characteristics—is used as both the key and the value. This allows $\psi$ to generate $\Delta_{ij}$ in a modality- and structure-aware manner, conditioned on the local pairwise discrepancy and auxiliary context. Rather than computing explicit gradient steps, $\psi$ learns to output increments that implicitly reflect the local descent direction. Since $\Delta_{ij}$ is embedded directly in the InfoNCE loss, the network is trained end-to-end via backpropagation, allowing it to align its output with the true coupled gradient field. As a result, $\psi$ captures both structure-aware priors and semantic context, producing informed corrections that reduce optimization tension and improve alignment robustness across diverse retrieval scenarios. We now consider the case where $c=v_{j}$ and the increment $\Delta_{ij}$ is added to the text side. In this setting, the two variables $t_{i}$ and $\Delta_{ij}$ share the same gradient flow:
$\nabla_{t_{i}}\mathcal{L}_{i}=\sum_{j=1}^{B}\nabla_{t_{\Delta_{ij}}}\mathcal{L}_{i},\qquad\nabla_{\Delta_{ij}}\mathcal{L}_{i}=\nabla_{t_{\Delta_{ij}}}\mathcal{L}_{i},$ (10)
where the gradient with respect to each perturbed anchor is given by:
$\nabla_{t_{\Delta_{ij}}}\mathcal{L}_{i}=(p_{ij}-y_{ij})\cdot\left[\frac{v_{j}}{\|t_{\Delta_{ij}}\|\cdot\|v_{j}\|}-\cos(t_{\Delta_{ij}},v_{j})\cdot\frac{t_{\Delta_{ij}}}{\|t_{\Delta_{ij}}\|^{2}}\right].$ (11)
@@
This formulation introduces a gradient redistribution effect: unlike standard InfoNCE where all gradients act directly on the shared anchor $t_{i}$, our pair-specific design allows each negative video $v_{j}$ to influence only its associated increment $\Delta_{ij}$. The positive pair $(t_{i},v_{i})$ contributes attraction gradients to both $t_{i}$ and $\Delta_{ii}$, facilitating alignment; meanwhile, each negative pair $(t_{i},v_{j}),j\neq i$, applies repulsion primarily to $\Delta_{ij}$. This leads to two key benefits: 1) tension relief — $\Delta_{ii}$ can absorb the alignment shift toward $v_{i}$, reducing the burden on $t_{i}$ to resolve large modality gaps in a single step; 2) false-negative suppression — repulsion from semantically similar negatives is redirected into their respective increments, preserving the semantic stability of the anchor representation. We apply this strategy under a symmetric InfoNCE loss:
$\mathcal{L}_{\text{info}}=-\frac{1}{2}\left(\frac{1}{B}\sum_{i=1}^{B}\log\frac{e^{s_{ii}/\tau}}{\sum_{j=1}^{B}e^{s_{ij}/\tau}}+\frac{1}{B}\sum_{j=1}^{B}\log\frac{e^{s_{ii}/\tau}}{\sum_{i=1}^{B}e^{s_{ij}/\tau}}\right),$ (12)
@@
Norm-Based Regularization of Trust-Region Radii
Each pair-specific increment $\Delta_{ij}$, predicted by the network $\psi$, approximates the optimal update direction under the first-order Taylor approximation. Its norm $\varepsilon_{ij}=\|\Delta_{ij}\|_{2}$ thus serves as a learned trust-region radius, regulating the extent of deviation that $t_{i}$ can take—regardless of direction—in order to absorb local alignment tension and semantic adjustment. To encourage structure in these radii, we regularize the intra-anchor variance of $\varepsilon_{ij}$ across each text anchor. Intuitively, semantically close pairs should have smaller $\varepsilon_{ij}$, while hard negatives benefit from larger ones—yet such structure must be inferred implicitly. We therefore promote norm diversity by minimizing the expected negative variance across anchors:
$\mathcal{L}_{\varepsilon}=\mathbb{E}_{t_{i}\sim\mathcal{B}_{t}}\left[-\text{Var}\left(\{\varepsilon_{ij}\}_{j=1}^{B}\right)\right],$ (13)
where $\mathcal{B}_{t}$ denotes the batch of text anchors. This encourages sharper separation between low-confidence and high-confidence pairs, without relying on external supervision. To prevent instability, we apply a lower bound:
$\mathcal{L}_{\varepsilon}=\max(\mathcal{L}_{\varepsilon},-\lambda),\quad\lambda>0.$ (14)
This simple yet effective regularization sharpens the implicit trust-region structure learned by $\psi$, guiding $\Delta_{ij}$ to reflect per-pair semantic variability while maintaining stable training dynamics. In the following section, we further complement this with a directional diversity loss to explicitly enhance the angular resolution of $\Delta_{ij}$ under each anchor.
Directional Diversity Regularization
@@
To enhance the expressiveness of the learned increments $\Delta_{ij}$, we introduce a directional regularization that encourages the directions of $\{\Delta_{ij}\}_{j=1}^{B}$ under each anchor $t_{i}$ to be diverse. This helps the model assign distinct update directions to different candidate videos, improving representational resolution and mitigating mode collapse.
We normalize each increment to obtain unit vectors $z_{ij}=\frac{\Delta_{ij}}{\|\Delta_{ij}\|_{2}}$ and define the regularization loss as the expected angular similarity across all anchor-specific direction sets:
@@
$\mathcal{L}_{\text{dir}}=\mathbb{E}_{t_{i}\sim\mathcal{B}_{t}}\left[\log\mathbb{E}_{j,k}\left[\exp\left(-\sigma\cdot(1-\langle z_{ij},z_{ik}\rangle)\right)\right]\right],$ (15)
@@
where $\sigma$ is a scale factor to control uniformity. This loss softly penalizes directional concentration, while still allowing nearby directions for semantically similar negatives—preserving flexibility under uncertainty. Combined with the norm-based regularization in Section 3.2.1, this term enables fine-grained control over both the magnitude and direction of each increment $\Delta_{ij}$, leading to more stable and structure-aware optimization.
KL Regularization via Information Bottleneck
To regularize the distribution of increments $\Delta_{ij}$ across text inputs for each video anchor $v_{j}$, we introduce a Information Bottleneck (IB) [1, 35] via a KL divergence loss that aligns the empirical distribution with a standard normal prior:
@@
$\mathcal{L}_{\text{IB}}=\mathbb{E}_{v_{j}\sim\mathcal{B}_{v}}\left[\mathrm{KL}\left(\mathcal{N}(\mu_{j},\sigma_{j}^{2})\,\|\,\mathcal{N}(0,I)\right)\right],$ (16)
where $\mu_{j}$ and $\sigma_{j}^{2}$ are the mean and variance of $\{\Delta_{ij}\}_{i=1}^{B}$ over the text batch for each $v_{j}$. This loss can be interpreted through the lens of the Information Bottleneck principle: it encourages $\Delta$ to retain minimal information about $v_{j}$ beyond what is necessary for pairwise alignment. By bounding the mutual information $I(\Delta;v_{j})$, the model avoids encoding redundant video-specific patterns in $\Delta$. This regularization complements the norm and direction constraints by promoting centered, isotropic, and information-efficient corrections, further enhancing generalization and stability.
Although the input configuration of $\psi$ may vary across datasets, we consistently apply the KL regularization along the video dimension, as video-side representations typically exhibit higher redundancy and local similarity. Regularizing the set of increments associated with each video anchor helps suppress noise and stabilizes learning by enforcing consistent structure across semantically similar pairs. See Appendix.B for an ablation study.
@@ 4 Experiment
Datasets, Metrics and Implementation Details
Datasets. We evaluate our method on four standard text-video retrieval benchmarks: MSR-VTT [45], DiDeMo [2], MSVD [6], and ActivityNet Captions [24]. MSR-VTT contains 10K videos with 20 captions each; we follow the 1K-A test split. DiDeMo includes 10K videos segmented into 5-second clips, each annotated with multiple sentences. MSVD consists of 1.9K short video clips with English captions. ActivityNet Captions provides dense annotations for 20K long-form videos with multiple temporally grounded descriptions.
Metrics. We choose Recall at rank K={1, 5, 10} (R@K), Median Rank (MdR), and mean rank (MnR) to evaluate the retrieval performance.
Implementation Details. We adopt CLIP (ViT-B/32) [33] as the base bi-encoder, equipped with a 4-layer Temporal Transformer [36] for video encoding. Following prior works [29, 15, 26, 39], we use 32-word captions and 12 video frames for MSR-VTT and MSVD, and 64-word captions with 64 frames for DiDeMo and ActivityNet Captions due to their longer video durations. We use the Adam optimizer [12] with linear warm-up, as in prior works. The learning rate is set to $1\text{e}{-7}$ for CLIP’s text and visual encoders, and $1\text{e}{-4}$ for all other modules (except $3\text{e}{-4}$ for ActivityNet). We set $\tau=0.01$, $\sigma=2$, and $\lambda=0.5$. All experiments use a batch size of 128. We train the model for 5 epochs on MSR-VTT, MSVD, and DiDeMo, and 10 epochs on ActivityNet Captions. All experiments are conducted on 4 to 8 GPUs including RTX 4090, A100 and V100.
Table 1: Comparison results on MSRVTT dataset on Text-to-Video Retrieval and Video-to-Text Retrieval.
@@
CLIP4Clip [29] 44.5 71.4 81.6 2.0 15.3 42.7 70.9 80.6 2.0 11.6
X-Pool [15] 46.9 72.8 82.2 2.0 14.3 44.4 73.3 84.0 2.0 9.0
TS2-Net [28] 47.0 74.5 83.8 2.0 13.0 45.3 74.1 83.7 2.0 9.2
EMCL-Net [19] 46.8 73.1 83.1 2.0 12.8 46.5 73.5 83.5 2.0 8.8
UATVR [14] 47.5 73.9 83.5 2.0 12.3 46.9 73.8 83.8 2.0 8.6
ProST [26] 48.2 74.6 83.4 2.0 12.4 46.3 74.2 83.2 2.0 8.7
HBI [20] 48.6 74.6 83.4 2.0 12.0 46.8 74.3 84.3 2.0 8.9
Diffusion [22] 49.0 75.2 82.7 2.0 12.1 47.7 73.8 84.5 2.0 8.8
EERCF [34] 47.8 74.1 84.1 - - 44.7 74.2 83.9 - -
MPT [49] 48.3 72.0 81.7 - 14.9 46.5 74.1 82.6 - 11.8
@@
Table 2: Comparison results on DiDeMo, ActivityNet Captions, and MSVD datasets on Text-to-Video Retrieval.
@@
Comparison with Other Methods
Table 1 and Table 2 shows the performance of our method across four standard text-video retrieval benchmarks. As seen, our approach consistently outperforms recent state-of-the-art methods on MST-VTT, ActivityNet and DiDeMo, and a comparable results in MSVD.
Ablative Analysis
Table 3: Ablation Studies on Modality Choice and Loss Combinations. left: We perform Text-to-Video Task on three datasets on different context modality type of our module $\psi$. right: We perform Text-to-Video Task on MSR-VTT for all losses ablation. First row refers to our baseline.
Context Modality Choice of $\psi$
Modality Dataset R@1↑ R@5↑ R@10↑ MnR↓
$c=t_{i}$ MSRVTT 47.4 73.5 82.1 12.9
$c=v_{j}$ MSRVTT 49.1 73.3 82.2 12.4
$c=t_{i}$ ActivityNet 42.6 73.6 84.4 6.8
$c=v_{j}$ ActivityNet 40.2 72.2 83.6 8.1
$c=t_{i}$ DiDeMo 46.5 74.3 82.6 12.3
$c=v_{j}$ DiDeMo 47.6 75.4 83.1 12.0
Losses Ablation

### v5

@@ 2.2 Gap-Aware Increment Modeling via Pair-Specific $\Delta_{ij}$
$\Delta_{ij}^{(t)}=\psi\left(\eta(v_{j}-t_{i}),\mathbf{C}\,;\,\Theta^{(t)}\right),$ (7)
where $\eta\in\{-1,+1\}$ controls whether the increment encodes video-specific or text-specific residual semantics. Here, $t_{i}$ denotes the [CLS] embedding of the text sequence, and $v_{j}$ is the mean-pooled representation of the video frame features $\mathbf{V}_{\text{frame}}$. We implement $\psi$ as a Cross-Attention module, with the semantic gap $v_{j}-t_{i}$ as the Query and the feature sequence $\mathbf{V}_{\text{frame}}\in\mathbb{R}^{N_{v}\times D}$ (or $\mathbf{T}_{\text{word}}\in\mathbb{R}^{N_{w}\times D}$) as the context $\mathbf{C}$ (i.e., the Key and Value). Intuitively, the query encodes what semantics are present in $v_{j}$ but missing in $t_{i}$ (or vice versa); sequence features carrying such semantics are assigned higher weights, so that the aggregated increment $\Delta_{ij}$ acts as a semantic patch correcting $t_{i}$. Rather than computing explicit descent steps, $\psi$ learns to output increments that approximate the coupled descent direction while incorporating structure-aware priors from the context. Since $\Delta_{ij}$ directly enters the InfoNCE loss, $\psi$ is optimized end-to-end via backpropagation, enabling its outputs to align with the true gradient field. We now consider the case where $\mathbf{C}=\mathbf{V}_{\text{frame}}$ and the increment $\Delta_{ij}$ is added to the text side. In this setting, the two variables $t_{i}$ and $\Delta_{ij}$ share the same gradient flow: $\nabla_{t_{i}}\mathcal{L}_{i}=\sum_{j=1}^{B}\nabla_{t_{\Delta_{ij}}}\mathcal{L}_{i}$ and $\nabla_{\Delta_{ij}}\mathcal{L}_{i}=\nabla_{t_{\Delta_{ij}}}\mathcal{L}_{i}$, where the gradient with respect to each perturbed anchor is given by:
$\nabla_{t_{\Delta_{ij}}}\mathcal{L}_{i}=\frac{(p_{ij}-y_{ij})}{\tau}\cdot\left[\frac{v_{j}}{\|t_{\Delta_{ij}}\|\cdot\|v_{j}\|}-\cos(t_{\Delta_{ij}},v_{j})\cdot\frac{t_{\Delta_{ij}}}{\|t_{\Delta_{ij}}\|^{2}}\right].$ (8)
@@
This formulation introduces a gradient redistribution effect: unlike standard InfoNCE where all gradients act directly on the shared anchor $t_{i}$, our pair-specific design allows each negative video $v_{j}$ to influence only its associated increment $\Delta_{ij}$. The positive pair $(t_{i},v_{i})$ contributes attraction gradients to both $t_{i}$ and $\Delta_{ii}$, facilitating alignment; meanwhile, each negative pair $(t_{i},v_{j}),j\neq i$, applies repulsion primarily to $\Delta_{ij}$. This leads to two key benefits: 1) tension relief: $\Delta_{ij}$ can absorb the gradient from $(t_{i},v_{j})$, reducing the burden on $t_{i}$ to decrease loss in a single step; 2) false-negative suppression: repulsion from semantically similar negatives is redirected into their respective increments, reducing the semantic bias of the anchor representation. We apply this strategy under a symmetric InfoNCE loss:
$\mathcal{L}_{\text{info}}=-\frac{1}{2}\left(\frac{1}{B}\sum_{i=1}^{B}\log\frac{e^{s_{ii}/\tau}}{\sum_{j=1}^{B}e^{s_{ij}/\tau}}+\frac{1}{B}\sum_{j=1}^{B}\log\frac{e^{s_{jj}/\tau}}{\sum_{i=1}^{B}e^{s_{ij}/\tau}}\right),$ (9)
@@
Norm-Based Regularization of Trust-Region Radii.
In our formulation, the increment $\Delta_{ij}$ predicted by $\psi$ is directly constrained within a trust region, thus its norm $\varepsilon_{ij}=\|\Delta_{ij}\|_{2}$ serves as the trust-region radius. This radius controls how far the corrected representation $t_{\Delta_{ij}}$ is allowed to deviate from $t_{i}$ in order to release optimization tension and adjust semantics. Intuitively, semantically similar pairs should yield smaller radii, while dissimilar pairs should allow larger ones. To encourage such structured variability, we regularize the intra-anchor distribution of radii by promoting norm diversity:
$\mathcal{L}_{\varepsilon}=-\mathbb{E}_{t_{i}\sim\mathcal{B}_{t}}\left[\text{Var}\left(\{\varepsilon_{ij}\}_{j=1}^{B}\right)\right],$ (10)
where $\mathcal{B}_{t}$ denotes the batch of text anchors. A lower bound $\max(\mathcal{L}_{\varepsilon},-\lambda)$ with $\lambda>0$ is applied to prevent instability. This regularization sharpens the implicit trust-region structure learned by $\psi$, guiding $\Delta_{ij}$ to reflect pairwise semantic variability in a stable manner.
Directional Diversity Regularization.
@@
To enhance the expressiveness of the learned increments $\Delta_{ij}$, we introduce a directional regularization that encourages the directions of $\{\Delta_{ij}\}_{j=1}^{B}$ under each anchor $t_{i}$ to be diverse. This helps the model assign distinct update directions to different candidate videos, improving the generalization of representations and mitigating mode collapse. Specifically, we normalize each increment to obtain unit vectors $z_{ij}=\frac{\Delta_{ij}}{\|\Delta_{ij}\|_{2}}$ and define the regularization loss as the expected angular similarity across all anchor-specific direction sets:
@@
$\mathcal{L}_{\text{dir}}=\mathbb{E}_{t_{i}\sim\mathcal{B}_{t}}\Big[\log\mathbb{E}_{j,k}\left[\exp\left(-\alpha\cdot(1-\langle z_{ij},z_{ik}\rangle)\right)\right]\Big],$ (11)
@@
where $\alpha$ is a scale factor to control uniformity. This loss softly penalizes directional concentration, while still allowing nearby directions for semantically similar negatives—preserving flexibility under uncertainty. Combined with the norm-based regularization, this term enables fine-grained control over both the magnitude and direction of each increment $\Delta_{ij}$, leading to more stable and structure-aware optimization.
2.3 Variational Information Bottleneck (VIB) for Semantic Increments
We motivate our regularization from the Information Bottleneck (IB) principle tishby2000information ; alemi2016deep , which seeks to maximize predictive information while suppressing nuisance factors. In our case, the increment $\Delta$ is optimized only by the gradient from its paired sample $(t_{i},v_{j})$ and thus lacks contrastive behavior, often collapsing into trivial solutions. We therefore treat $\Delta$ as an information bottleneck that extracts semantic signals from $(t_{i},v_{j})$ while discarding noise, effectively constraining it within a prior structure before allowing it to reduce the InfoNCE objective. Formally, this corresponds to maximizing the mutual information between $\Delta$ and the target semantics while minimizing its dependence on the input pair:
$\max_{p(\Delta|X)}I(\Delta;Y)-\beta\,I(\Delta;X),$ (12)
where $X=(t_{i},v_{j})$ denotes a input pair, $Y$ the label indicating whether it is a positive match, $\Delta$ the pair-specific increment (i.e, the latent variable $Z$), and $\beta$ trade-off between task term $I(\Delta;Y)$ and compression term $I(\Delta;X)$. Following the standard variational derivation (details in Appendix D), we obtain the objective
$\mathcal{L}_{\text{VIB}}:=\underbrace{-\mathbb{E}_{(t,v,y)}\mathbb{E}_{\Delta\sim q_{\psi}(\Delta|t,v)}\big[\log q_{\theta}(y|\Delta)\big]}_{\text{contrastive term }\mathcal{L}_{\text{info}}}+\beta\cdot\underbrace{\mathbb{E}_{(t,v)}\big[\mathrm{KL}\big(q_{\psi}(\Delta|t,v)\|r(\Delta)\big)\big]}_{\text{compression term }\mathcal{L}_{\text{IB}}},$ (13)
where $q_{\psi}(\Delta|t,v)$ denotes the variational encoder that parameterizes the posterior of $\Delta$ given $(t,v)$, $q_{\theta}(y|\Delta)$ is the variational classifier instantiated as a softmax predictor, and $r(\Delta)=\mathcal{N}(0,I)$ is the Gaussian prior used in the upper-bound regularization of the latent space.
In practice, we adopt a deterministic instantiation where the variational distribution $q_{\psi}(\Delta|t,v)$ collapses to a Dirac posterior centered at the network output $\Delta_{ij}$ (i.e., $\mu=\Delta_{ij}$), without uncertainty-aware sampling. Since the Dirac measure is singular with respect to any continuous prior, the KL divergence is ill-defined. To obtain a tractable and stable regularizer, we aggregate posteriors along the text dimension, leveraging the asymmetric nature of video–text data (i.e., videos typically contain higher redundancy and correspond to multiple semantically related texts). By the convexity of $\mathrm{KL}(\cdot\|r)$ and Jensen’s inequality, this yields a relaxation that also circumvents the singularity of the deterministic posterior:
$\mathbb{E}_{(t,v)}\big[\mathrm{KL}\left(q_{\psi}(\Delta|t,v)\|r(\Delta)\right)\big]$ $=\mathbb{E}_{v}\mathbb{E}_{t|v}\big[\mathrm{KL}\left(q_{\psi}(\Delta|t,v)\|r(\Delta)\right)\big]$ (14)
$\geq\mathbb{E}_{v}\big[\mathrm{KL}\big(\bar{q}_{\psi}(\Delta|v)\|r(\Delta)\big)\big],$
where $\bar{q}_{\psi}(\Delta|v):=\mathbb{E}_{t|v}[\,q_{\psi}(\Delta|t,v)\,]$ is the aggregated increment posterior for video $v$. This relaxation preserves the information-bottleneck effect while reducing sensitivity to text-side variability. The precise relationship between this relaxed KL term and the original mutual information $I(\Delta;X)$ is derived in Appendix E. Putting the relaxation into practice, we approximate $\bar{q}_{\psi}(\Delta|v_{j})$ with a Gaussian fitted to the set of increments $\{\Delta_{ij}\}_{i=1}^{B}$ associated with each video anchor $v_{j}$. This yields the following relaxed compression loss:
@@ 2.3 Variational Information Bottleneck (VIB) for Semantic Increments
$\mathcal{L}_{\text{IB}}^{\text{relax}}=\mathbb{E}_{v_{j}\sim\mathcal{B}v}\left[\mathrm{KL}\!\left(\mathcal{N}(\mu_{j},\sigma_{j}^{2})\|\mathcal{N}(0,I)\right)\right],$ (15)
where $\mu_{j}$ and $\sigma_{j}^{2}$ denote the mean and variance of increments $\{\Delta_{ij}\}_{i=1}^{B}$ over the text batch for each video $v_{j}$. This relaxed KL term operationalizes the bottleneck by regularizing increments at the video level, enforcing centered and isotropic corrections while avoiding over-penalization of text-side variability. We provide the overall training objective and inference details in Appendix F.
@@ 3 Experiment
Table 1: Comparison results on MSR-VTT dataset on Text-to-Video Retrieval and Video-to-Text Retrieval. DiCoSA jin2023text utilizes QB-Norm bogolin2022cross for inference and is grayed out for a fair comparison. Note that T2VLA wang2021t2vlad is a non-CLIP method.
@@
T2VLA wang2021t2vlad CVPR21 29.5 59.0 70.1 4.0 - 31.8 60.0 71.1 3.0 -
CLIP4Clip luo2022clip4clip Neurocomputing22 44.5 71.4 81.6 2.0 15.3 42.7 70.9 80.6 2.0 11.6
X-Pool gorti2022x CVPR22 46.9 72.8 82.2 2.0 14.3 44.4 73.3 84.0 2.0 9.0
TS2-Net liu2022ts2 ECCV22 47.0 74.5 83.8 2.0 13.0 45.3 74.1 83.7 2.0 9.2
EMCL-Net jin2022expectation NeurIPS22 46.8 73.1 83.1 2.0 12.8 46.5 73.5 83.5 2.0 8.8
UATVR fang2023uatvr ICCV23 47.5 73.9 83.5 2.0 12.3 46.9 73.8 83.8 2.0 8.6
DiCoSA jin2023text IJCAI23 47.5 74.7 83.8 2.0 13.2 46.7 75.2 84.3 2.0 8.9
ProST li2023progressive ICCV23 48.2 74.6 83.4 2.0 12.4 46.3 74.2 83.2 2.0 8.7
HBI jin2023video CVPR23 48.6 74.6 83.4 2.0 12.0 46.8 74.3 84.3 2.0 8.9
DiffusionRet jin2023diffusionret ICCV23 49.0 75.2 82.7 2.0 12.1 47.7 73.8 84.5 2.0 8.8
EERCF tian2024towards AAAI24 47.8 74.1 84.1 - - 44.7 74.2 83.9 - -
MPT zhang2024mpt ACM MM24 48.3 72.0 81.7 - 14.9 46.5 74.1 82.6 - 11.8
@@
Table 2: Comparison results on DiDeMo, ActivityNet Captions, and MSVD datasets on Text-to-Video Retrieval. Note that FROZEN bain2021frozen is a non-CLIP method.
@@
3.1 Experiment settings
Datasets and Metrics.
We evaluate our method on four standard text-video retrieval benchmarks: MSR-VTT xu2016msr , DiDeMo anne2017localizing , MSVD chen2011collecting , and ActivityNet Captions krishna2017dense . MSR-VTT contains 10K videos with 20 captions each; we follow the 1K-A validation split. DiDeMo includes 10K videos segmented into 5-second clips, each annotated with multiple sentences. MSVD consists of 1.9K short video clips with English captions. ActivityNet Captions provides dense annotations for 20K long-form videos with multiple temporally grounded descriptions. We choose Recall at rank K={1, 5, 10} (R@K), Median Rank (MdR), and Mean Rank (MnR) to evaluate the retrieval performance.
Implementation Details.
We adopt CLIP (ViT-B/32) radford2021learning as the base dual-encoder, equipped with a 4-layer Temporal Transformer vaswani2017attention following the CLIP vision encoder for video encoding. Following prior works luo2022clip4clip ; gorti2022x ; li2023progressive ; DRLTVR2022 , we use 32-word captions and 12 video frames for MSR-VTT and MSVD, and 64-word captions with 64 frames for DiDeMo and ActivityNet Captions due to their longer video durations. We use the Adam optimizer diederik2014adam with linear warm-up, as in prior works. The learning rate is set to $1\text{e}{-7}$ for CLIP’s text and visual encoders, and $1\text{e}{-4}$ for all other modules. We set $\beta=0.07$, $\tau=0.01$, $\alpha=2$, and $\lambda=0.5$ for MSR-VTT. All experiments use a batch size of 128. We train the model for 5 epochs on MSR-VTT, MSVD, and DiDeMo, and 10 epochs on ActivityNet Captions. All experiments are conducted on 4 to 8 GPUs including RTX 4090, A100 and V100.
3.2 Comparison with Other Methods
Table 1 and Table 2 shows the performance of our method across four standard text-video retrieval benchmarks. As seen, our approach consistently outperforms recent state-of-the-art methods on MSR-VTT, ActivityNet, DiDeMo and MSVD.
3.3 Ablative Analysis
All ablation studies are conducted on the MSR-VTT 1k-A validation set. In addition to the ablations presented below, we further provide results on the lower bound coefficient $\lambda$ of $\mathcal{L}_{\varepsilon}$, the interaction between the context modality type $\mathbf{C}$ of $\psi$ and the direction indicator $\eta$, the scale factor $\alpha$ of $\mathcal{L}_{\text{dir}}$, the choice of anchor in $\mathcal{L}_{\text{IB}}^{\text{relax}}$ and hard negative methods comparison in Appendix C.
Table 3: Ablation on losses combination on Text-to-Video Retrieval results on MSR-VTT 1k-A. First row denotes the baseline.